# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

### v3

@@ Abstract
The launch of Grokipedia—an AI-generated encyclopedia developed by Elon Musk’s xAI—was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries using the Grok large language model. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study conducts a large-scale computational comparison of more than 17,000 matched article pairs from the 20,000 most-edited English Wikipedia pages. Using metrics spanning lexical richness, readability, reference density, structural features, and semantic similarity, we assess how closely the two platforms align in form and substance. We find that Grokipedia articles are substantially longer and contain significantly fewer references per word. Moreover, Grokipedia’s content divides into two distinct groups: one that remains semantically and stylistically aligned with Wikipedia, and another that diverges sharply. Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited news sources, concentrated primarily in entries related to politics, history, and religion. These findings suggest that AI-generated encyclopedic content diverges from established editorial norms—favouring narrative expansion over citation-based verification. The implications highlight emerging tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
@@ 1 Introduction
In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia. [^2] See [GrokipediaWP, AP2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025]. According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[GrokipediaWP, BI2025, AP2025] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[Verge2025, Wired2025, Guardian2025, Time2025]
@@
Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We address this by pairing articles across platforms and comparing them along three similarity dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), and (3) stylistic similarity (readability, lexical diversity, part-of-speech composition). In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
We also measure the political bias of each article by aggregating the bias scores of its cited references (see Methods for details). To contextualize these patterns, we classify all articles into topical domains and then analyse how both similarity and the shift in political bias vary across topics.
To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[yasseri2012dynamics] Among these, approximately 17,700 titles had valid, nontrivial matches on Grokipedia.
This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure. Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI. [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, GrogConversation2025].
@@ 2.1 Sampling and Data Collection
We analyzed the 20,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts. Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [yasseri2012dynamics]. To ensure that our comparison focused on substantive articles, we excluded all list-style pages as well as titles that were date- or year-like rather than topical. Specifically, we removed (a) pages with the title ”List of Topic” (e.g., UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
@@
HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11) with polite delays and a standard user-agent header, and then parsed using BeautifulSoup4. After host-aware text extraction (described below), we retained only article pairs in which both platforms produced at least 500 words of clean prose. Of the original 20,000 target titles, 17,790 matched pairs met these criteria and formed the final analytical sample.
2.2 Text Cleaning and Feature Extraction
We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximise content fidelity.
For Wikipedia, extraction was restricted to the #mw-content-text container and limited strictly to <p> and <li> elements. Elements such as infoboxes, hatnotes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
For Grokipedia, a more adaptive extractor identified the main article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high link-density regions. Scripts, styles, and sidebars were removed in both cases.
Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, and stylistic metrics was computed.
  • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
  • Lexical density and reading time: ratio of unique to total words and estimated reading time assuming 200 words per minute.
@@ 2.3 Similarity Measures Between Platforms
To quantify cross-platform alignment, we computed a suite of eight similarity measures grouped into four conceptual domains:
@@
    Lexical similarity: cosine similarity of TF–IDF vectors (1–2 grams, English stop-words removed) and unigram Jaccard index.
@@
    N-gram overlap: overlap coefficients for 1-, 2-, and 3-gram sequences, capturing local phrase reuse.
@@
    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[reimers2019sentence], and contextual similarity via BERTScore F1 for the first 50 sentences per article [zhang2019bertscore].
@@
    stylistic similarity: The Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
@@ 2.4 Political Bias ratings
Outlet-level political-leaning scores were taken from the News Media Bias and Factuality dataset [sanchez2024mapping], mapped to a symmetric numeric scale, and used only when a numeric bias was available for a domain or its brand variant (e.g., bbc.com/bbc.co.uk). Non-news and reference/utility hosts (e.g., doi.org, archive.org, worldcat.org, major platforms, and *.gov/*.edu) were excluded. For brand merging we used eTLD+1 normalization plus a brand key (leftmost label).
@@
For each article, bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift). Centers of mass in Figure 6 were derived using Gaussian kernel density estimation (KDE) applied to the raw article-level points, computed separately for the low- and high-similarity partitions. For the pie charts in Figure 5, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 40% cumulative share of referenced news outlets. Slice colors encode outlet bias, and panel titles report the mention-weighted average platform bias normalized to ±100%.
2.5 Topics
@@
To categorize the topics of the articles, we developed a multi-stage large-language-model (LLM) classification pipeline. We first constructed a fixed taxonomy of fourteen semantically distinct categories (e.g., History, Politics and Political Figures, Science and Technology, Literature and Art). Article titles were processed in batches of up to 300 and passed to the GPT-5-nano-2025-08-07 model with strict instructions to assign each title to exactly one category and to avoid the fallback use of “Other” whenever a reasonable category was available. The model returned a JSON-formatted list of titles grouped by category, and the script automatically retried up to five times if the output contained formatting errors. All valid batch outputs were then combined into an initial classification table mapping each title to its assigned category.
@@
Because LLM-generated labels can occasionally be inconsistent across runs, we used a multi-round refinement procedure to improve reliability. The entire classification process was repeated across several independent rounds using the same model and prompt. After each round, titles labeled “Other” were separated and given to the model for reclassification. We then merged the rounds. Any article still lacking a category after all rounds was assigned “Other”. This multi-round integration approach increased coverage, reduced noise, and yielded a more stable and robust set of topic labels for all article pairs.
Articles were assigned to one of the following categories:
  • History
  • Politics and Political Figures
  • Geography, Nations, and Cities
  • Music, Movies, and Entertainment Celebrities
  • Sports, Athletes, and Teams
  • Science, Technology, and Knowledge
  • Religion and Ideologies
  • Brands and Products
  • Languages
  • Literature and Art
  • Health and Environment
  • Business and Infrastructure
  • Animals and Nature
  • Other
@@ 2.6 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. Descriptive statistics (mean $\pm$ SD) were calculated for all features. Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering. Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib.
@@ 2.7 Limitations
The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia. Because edit frequency correlates with controversy and visibility [yasseri2012dynamics], the sample likely overrepresents politically or culturally salient topics. In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose. Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
@@ 3.5 Political Bias
Figure 6: Bias shift vs. article similarity. Each point represents an article pair. The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions (higher values indicate greater similarity), and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia). Positive values indicate a shift toward right-leaning sources, while negative values indicate a shift toward left-leaning sources. Colored markers denote the centres of mass for the low-similarity and high-similarity groups. As similarity increases, bias shifts converge toward zero; when similarity is low, the magnitude and dispersion of bias shifts increase.
To further examine differences in the sources cited within the least similar article pairs, we focus on the subset of articles with a negative combined similarity score. Figure 5 displays the top 40% most frequently cited news sources within this low-similarity group for each platform. Each wedge is colour-coded according to the ideological bias of the outlet. As shown in the figure, the weighted average bias score of Grokipedia articles in this subset is approximately 25% left-leaning, compared with 31% left-leaning for Wikipedia. This difference indicates a relative rightward shift in the sourcing of the least similar Grokipedia articles, consistent with the pattern observed in the bias–similarity relationship above.
Figure 7: Top news sources among low-similarity articles (negative combined similarity score). Wedges show each outlet’s share of all mentions within this subset; colors reflect the outlet’s political-leaning score from the public News Media Bias and Factuality dataset. [^5] https://huggingface.co/datasets/sergioburdisso/news_media_bias_and_factuality Titles report the overall (news-only) mention-weighted bias for each platform (normalized to $\pm 100\%$).

### v4

@@ Abstract
The launch of Grokipedia—an AI-generated encyclopedia developed by Elon Musk’s xAI—was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries using the Grok large language model. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study conducts a large-scale computational comparison of 17,790 matched article pairs from the 20,000 most-edited English Wikipedia pages. Using metrics spanning lexical richness, readability, reference density, structural features, and semantic similarity, we assess how closely the two platforms align in form and substance. We find that Grokipedia articles are substantially longer and contain significantly fewer references per word. Moreover, Grokipedia’s content divides into two distinct groups: one that remains semantically and stylistically aligned with Wikipedia, and another that diverges sharply. Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited sources, concentrated primarily in entries related to politics, history, and religion. More broadly, the findings indicate that AI-generated encyclopedic content departs from established editorial norms, favoring narrative expansion over citation-based verification, raising questions about transparency, provenance, and the governance of knowledge in automated information systems.
@@ 1 Introduction
In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia. [^2] See [GrokipediaWP, apnews2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025]. According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance [GrokipediaWP, BI2025, apnews2025]. Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models [Verge2025, Wired2025, Guardian2025, Time2025].
@@
Against this backdrop, our study asks a concrete, testable question: for titles that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We address this by pairing articles across platforms and comparing them along three similarity dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), and (3) stylistic similarity (readability, lexical diversity, part-of-speech composition). In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
We also measure the political bias of each article by aggregating the bias scores of its cited sources (see Methods for details). To contextualize these patterns, we classify all articles into topical domains and then analyse how both similarity and the shift in political bias vary across topics.
To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience [yasseri2012dynamics]. Among these, 17,790 titles had valid, nontrivial matches on Grokipedia.
This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure. Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI  [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, yasseri2025conversation]..
@@ 2.1 Sampling and Data Collection
We analyzed the 20,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts. To ensure that our comparison focused on substantive articles, we excluded all list-style pages as well as titles that were date- or year-like rather than topical. Specifically, we removed (a) pages with the title ”List of Topic” (e.g., List of UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
@@
HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11), and then parsed using BeautifulSoup4. After host-aware text extraction (described below), we retained only article pairs in which both platforms produced at least 500 words of clean prose. Of the original 20,000 target titles, 17,790 matched pairs met these criteria and formed the final analytical sample.
2.1.1 Host-aware text and feature extraction
We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
For Wikipedia, extraction was restricted to the #mw-content-text container and limited to <p> and <li> elements. Infoboxes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
For Grokipedia, we used a more adaptive extractor to identify the primary article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high–link-density regions. Scripts, style elements, and sidebars were removed for both platforms.
We tokenized each cleaned article into sentences and words using nltk’s Punkt tokenizer.
From the extracted text of each article, we computed a set of stylistic and readability metrics, in addition to lexical density and estimated reading time. All features were calculated separately for each platform.
  • Average sentence length: Computed as the mean number of tokens per sentence in the extracted text. This measure captures syntactic complexity, with longer sentences generally reflecting more complex or formally structured writing.
  • Lexical diversity (type–token ratio): Defined as the number of unique word types divided by the total number of word tokens. Higher values indicate richer vocabulary usage and lower repetition within the text.
  • Flesch–Kincaid grade level: A widely used readability formula that estimates the U.S. school grade level required to understand the text. It combines average sentence length and average syllables per word, with higher scores indicating more difficult material.
  • Gunning–Fog index: A readability metric evaluating the number of complex words (three or more syllables) and sentence length. It approximates the years of formal education needed to comprehend a text on first reading and is commonly used to quantify formal, technical, or academic writing.
  • Part-of-speech (POS) composition: Using spaCy’s en_core_web_sm tagger, we computed the proportion of nouns, verbs, adjectives, and adverbs in each article. POS distribution provides a coarse-grained linguistic profile of stylistic variation, such as noun-dense factual exposition versus verb-heavy narrative text.
  • Lexical density: Calculated as the ratio of unique words (types) to total words (tokens). Lexical density reflects how information-dense a text is; higher values indicate a more compact and content-rich lexical profile.
  • Estimated reading time: Derived by dividing the total word count by a constant reading rate of 200 words per minute. This measure offers a standardized estimate of the amount of content a typical reader must process.
@@ 2.2 Similarity Measures Between Platforms
To quantify alignment between articles on the two platforms, we computed four classes of similarity measures, each capturing a distinct dimension of textual resemblance:
@@
    Lexical similarity: Lexical similarity was assessed using two complementary approaches. First, we computed cosine similarity between TF–IDF vectors constructed from 1–2 gram features with standard English stop-words removed. This measure captures similarity in word- and phrase-level vocabulary usage weighted by informativeness. Second, we calculated the unigram Jaccard index, defined as the ratio of the intersection to the union of unique word sets across platforms. This index evaluates the extent to which the two texts draw from the same vocabulary.
@@
    N-gram overlap: To capture local phrase reuse and short-range structural similarity, we computed overlap coefficients for 1-, 2-, and 3-gram sequences. Each coefficient is defined as the size of the intersection of n-grams divided by the smaller of the two n-gram sets, providing a symmetric measure of shared phrase content independent of text length.
@@
    Semantic similarity: We quantified semantic alignment using two embedding-based approaches. Cosine similarity between SentenceTransformer embeddings
    (all-MiniLM-L6-v2) measured the overall semantic closeness of the full articles in a dense representation space. In addition, we computed contextual similarity using BERTScore F1 on the first 50 sentences of each article, which evaluates token-level semantic correspondence using contextual embeddings [reimers2019sentencebert]. Together, these metrics assess similarity in meaning beyond surface-level wording.
@@
    Stylistic similarity: Stylistic similarity was evaluated by computing the Manhattan distance between the two articles’ stylistic feature vectors, which included sentence length, lexical diversity, readability metrics, and part-of-speech composition. The resulting distances were transformed linearly to a 0–1 scale, where higher values indicate stronger similarity. This measure captures convergence in writing style independent of content.
@@ 2.3 Political Bias ratings
We extracted all outbound hyperlinks contained in each article and mapped their host domains to political-leaning scores using the News Media Bias and Factuality dataset [sanchez2024mapping]. Scores from this dataset were converted to a symmetric numeric scale (-1:left to 1:right) and were included only when a quantitative bias value was available for the domain or any of its recognized brand variants (e.g., bbc.com/bbc.co.uk).
To consolidate hosts referring to the same outlet, we applied eTLD+1 normalization supplemented by a brand key derived from the leftmost label. This procedure ensured consistent attribution across domain variants and subdomain structures.
@@
The bias score of each article is the weighted average of the bias score of each domain based on the number of times the domain is mentioned in the page. For each article pair, the bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift). The centres of mass shown in Figure 1 were obtained via Gaussian kernel density estimation (KDE) applied to the article-level scatter points, computed separately for the low- and high-similarity partitions.
For the pie charts in Figure 1, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 5 of the cumulative share of referenced news outlets. In this analysis, we focused exclusively on news sources, removing reference or utility hosts (e.g., doi.org, archive.org, worldcat.org), major platforms, and government or educational domains (.gov, .edu).
2.4 Categorisation
@@
To assign each article to a topical domain, we implemented a multi-stage large-language-model (LLM) classification pipeline. Article titles were processed in batches of up to 300 and submitted to the gpt-5-nano-2025-08-07 model with explicit instructions to assign each title to exactly one topic and to avoid the use of “Other” unless no reasonable category applied. The model returned a JSON-formatted list of titles grouped by category. Outputs with formatting errors were automatically resubmitted, with up to five retries per batch. All valid outputs were merged into an initial table mapping each title to a category.
@@
Because LLM-generated labels can vary across runs, we conducted several independent classification rounds using the same model and prompt. After each round, titles labeled as “Other” were isolated and resubmitted for reclassification. The results were then merged across rounds. Titles that remained unclassified after all iterations were ultimately assigned to “Other”. This multi-round integration procedure increased coverage, reduced inconsistency, and yielded a more stable set of topic labels for all article pairs.
As a final validation step, we submitted each provisional category (batched in sets of 200 entries) to the gpt-4.1-mini model and instructed it to flag and propose revised labels for any titles that appeared misclassified. All suggested corrections were manually reviewed before inclusion in the final categorisation.
Articles were assigned to one of the following topical domains: history, politics, geography, entertainment, sports, science and technology, religion and ideologies, brands and products, languages, literature and art, health and environment, business and infrastructure, animals and nature, or other.
@@ 2.5 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. We calculated the descriptive statistics (mean $\pm$ SD) for all features. We evaluated Differences between Grokipedia and Wikipedia using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering. Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib.
@@ 2.6 Limitations
Our dataset is affected by several limitations. First, the dataset is drawn from Wikipedia’s 20,000 most-edited English-language pages, of which 17,790 have counterparts on Grokipedia. This selection likely overrepresents high-profile and contentious topics—those most prone to editing disputes and ideological scrutiny [yasseri2012dynamics]. Second, the similarity metrics employed—lexical, semantic, structural, and stylistic—capture textual form and alignment but do not assess factual accuracy or ideological framing. As a result, hallucinated claims, selective omissions, or more subtle rhetorical biases remain beyond the scope of automated comparison. Third, the bias measure is based on the ideological orientation of source domains rather than individual articles, relies on a limited dataset, and is grounded in the U.S. political context, which may not generalize to other media systems. Fourth, both platforms are dynamic: Grokipedia and Wikipedia continue to evolve, and our analysis reflects only a snapshot of their content at the time of data collection. Finally, Grokipedia’s underlying training data, retrieval mechanisms, and editorial interventions remain opaque, precluding full provenance auditing and limiting causal inference about the sources of observed bias.
@@ 3.5 Political Bias
Figure 6: Bias shift vs. article similarity. Each point represents an article pair. The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions, and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia). Positive values indicate a shift toward right-leaning sources, while negative values indicate a shift toward left-leaning sources. Colored markers denote the centres of mass for the low- and high-similarity groups. Inset: Pie charts show the top five news domains cited in Grokipedia (left) and Wikipedia (right) among articles with negative combined similarity scores, weighted by citation frequency. Slice areas represent relative contribution, and colors encode political bias.
The inset of Figure 6 displays the five most frequently cited news sources within this low-similarity group for each platform, with wedges color-coded by ideological bias. As shown, the least similar Grokipedia articles exhibit a relative rightward shift in sourcing; however, they remain predominantly left-leaning.